### v1

@@ 1 Introduction
where $h:\Re^{d}\rightarrow\Re$ is twice continuously differentiable convex function and is essentially strictly convex, $\mathcal{Q}:\Re^{m\times n}\rightarrow\Re^{d}$ and $\mathcal{B}:\Re^{m\times n}\rightarrow\Re^{l}$ are two linear operators, $C\in\Re^{m\times n}$ and $b\in\Re^{l}$ are given data and $\mathcal{P}\subseteq\Re^{l}$ is a given convex polyhedra cone. The function $\|X\|_{2}$ is the spectral norm function, namely, the largest singular value of $X$. Without loss of generality, we suppose that $m\leq n$ in what follows. Problem (1.1) has a wide range of applications in various fields, such as matrix approximation problems [1], matrix chebyshev polynomial problems [2] and $H_{\infty}$ synthesis problems [3, 4, 5], etc. Recently, spectral norm regularized problem has also been applied in deep learning and neural networks problems [6, 7, 8].
@@ 2 Notation and Preliminaries
Let $\mathcal{X}$ and $\mathcal{Y}$ be two finite dimensional real Euclidean spaces. Given $\rho>0$, define the ball $\mathbb{B}_{\rho}(z):=\{z^{\prime}\in\mathcal{X}~{}|~{}\|z^{\prime}-z\|\leq\rho\}$, $\forall z\in\mathcal{X}$. Let $D\subseteq\mathcal{X}$ be a non-empty closed convex set. For any $z\in D$, the tangent cone (see [17, Definition 6.1]) of $D$ at $z$ is defined by ${\mathcal{T}}_{D}(z):=\{d\in\mathcal{X}~{}|~{}\exists z^{k}\rightarrow z$ with $z^{k}\in D$ and $\tau^{k}\downarrow 0$ s.t. $(z^{k}-z)/\tau^{k}\rightarrow d\}$. We use $\mathcal{N}_{D}(x):=\{d\in\mathcal{X}:\langle d,z-x\rangle\leq 0~{}\forall z\in D\}$ to denote the normal cone of $D$ at $x\in D$. Denote by $\delta_{D}(x)$ the indicator function over $D$, i.e., $\delta_{D}(x)=0$ if $x\in D$, and $\delta_{D}(x)=\infty$ if $x\notin D$. Define the support function of the set $D$ as $\sigma(y,D):=\sup_{x\in D}\langle x,y\rangle,y\in\mathcal{X}$. For a given $x\in\mathcal{X}$, define $\Pi_{D}(x):=\arg\min\{\|d-x\|~{}|~{}d\in D\}$ as the projection mapping onto $D$. Suppose the function $f:\mathcal{X}\rightarrow(-\infty,+\infty]$ is proper closed convex, denote by ${\rm dom}f:=\{x\mid f(x)<\infty\}$ the effective domain of $f$, by $f^{*}$ the conjugate of $f$ and by $\partial f$ the subdifferential of $f$. For more details, one can refer standard convex analysis [17]. For any convex cone $K$, denote by $K^{\circ}:=\{x\in\Re^{l}\mid\langle x,z\rangle\leq 0,\forall z\in K\}$ the polar of $K$.
@@ 3 Variational Analysis of the Spectral Norm Function
In this section, we will provide a direct expression for the directional derivative of the proximal mapping of the spectral norm function. Then discuss the relationship between directional derivatives of the proximal mapping, the so called “sigma term” and the critical cones. In the following text, we always use $\theta$ to represent the spectral norm function, i.e., $\theta(\cdot)=\|\cdot\|_{2}$.
@@
    If $\lambda_{1}(S(\widetilde{H}_{\alpha\alpha}))\neq 0$, then $H\in(\mathcal{C}_{\theta}(P,W))^{\circ}\iff\varphi^{*}_{(W,H)}(P)=0$ and $H\in\mathcal{C}_{\theta^{*}}(W,P)$, where $\varphi^{*}_{(W,H)}(P)$ is the conjugate function of $\varphi_{(W,H)}(\cdot)=(\theta^{*})^{\prime\prime}(W;H,\cdot)$.
@@
    If $\lambda_{1}(S(\widetilde{H}_{\alpha\alpha}))=0$, then $H\in(\mathcal{C}^{0}_{\theta}(P,W))^{\circ}\Longleftarrow\varphi^{*}_{(W,H)}(P)=0$ and $H\in\mathcal{C}_{\theta^{*}}(W,P)$, where $\mathcal{C}^{0}_{\theta}(P,W)$ represents the critical cone $\mathcal{C}_{\theta}(P,W)$ in the case of $\tau=0$, i.e.,
  3. (iii)
    If $\lambda_{1}(S(\widetilde{H}_{\alpha\alpha}))\neq 0$, then $H\in(\mathcal{C}_{\theta^{*}}(P,W))^{\circ}\iff\psi^{*}_{(P,H)}(W)=0$ and $H\in\mathcal{C}_{\theta}(P,W)$.
  4. (iv)
    If $\lambda_{1}(S(\widetilde{H}_{\alpha\alpha}))=0$, then $H\in(\mathcal{C}_{\theta^{*}}(P,W))^{\circ}\Longleftarrow\psi^{*}_{(P,H)}(W)=0$ and $H\in\mathcal{C}^{0}_{\theta}(P,W)$.
@@
which means that $H\in(\mathcal{C}_{\theta}(P,W))^{\circ}$. When $\tau^{\prime}$ in $\mathcal{C}_{\theta}(P,W)$ is not equal to 00, i.e., $\lambda_{1}(S(\widetilde{D}_{\alpha\alpha}))\neq 0$, it is not hard to prove that vice versa. When $\tau^{\prime}$ in $\mathcal{C}_{\theta}(P,W)$ is equal to 00, we can not get ${\rm trace}(\widetilde{H}_{\alpha\alpha})=0$, i.e., the reverse conclusion is not true. Then the proof of part (i) and (ii ) are completed.
To prove (iii), it is easy to see that $\psi^{*}_{(P,H)}(W)=0$ and $H\in\mathcal{C}_{\theta}(P,W)$ hold if and only if there exists $\tau^{\prime\prime}\in\Re$ such that
@@
which can be proved to be equivalent to $H\in(\mathcal{C}_{\theta^{*}}(W,P))^{\circ}$ by using the same method as in part (i) when $\tau^{\prime\prime}\neq 0$. When $\tau^{\prime\prime}=0$, we also can not get the reverse conclusion. ∎
@@
Let $W\in\partial\theta(P)$ and $Q=P+W$ admit the SVD as in (3.1). Let the index sets $\alpha$, $\beta$, $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ and $c$ be defined as (3.2), (3.4) and (3.6). Given any $D\in\Re^{m\times n}$, denote $\widetilde{D}=U^{T}DV$. The function $\psi^{*}_{(P,D)}(W)$ is defined in (3.14). For any $D\in\Re^{m\times n}$, if $H={\rm Prox}^{\prime}_{\theta}(Q;H+D)$ holds, then $H\in\mathcal{C}^{0}_{\theta}(P,W)$ and $\langle H,D\rangle=-\psi^{*}_{(P,D)}(W)$.
@@
where $\Xi_{\alpha\alpha}$ is defined in (3.13). We first verify $H\in\mathcal{C}^{0}_{\theta}(P,W)$. The equality
@@
From part (i) of proposition 3.1, we obtain that $H\in\mathcal{C}^{0}_{\theta}(P,W)$.
@@
${\rm Prox}^{\prime}_{\theta}(Q;D)=0\Longleftrightarrow D\in(\mathcal{C}^{0}_{\theta}(P,W))^{\circ}.$

### v2

@@ 1 Introduction
where $h:\Re^{d}\rightarrow\Re$ is twice continuously differentiable convex function and is essentially strictly convex, $\mathcal{Q}:\Re^{m\times n}\rightarrow\Re^{d}$ and $\mathcal{B}:\Re^{m\times n}\rightarrow\Re^{l}$ are two linear operators, $C\in\Re^{m\times n}$ and $b\in\Re^{l}$ are given data and $\mathcal{P}\subseteq\Re^{l}$ is a given convex polyhedra cone. The function $\|X\|_{2}$ is the spectral norm function, namely, the largest singular value of $X$. Without loss of generality, we suppose that $m\leq n$ in what follows. For convenience, in the following text, we always use $\theta$ to represent the spectral norm function, i.e., $\theta(\cdot)=\|\cdot\|_{2}$. Problem (1.1) has a wide range of applications in various fields, such as matrix approximation problems [1], matrix chebyshev polynomial problems [2] and $H_{\infty}$ synthesis problems [3, 4, 5], etc. Recently, spectral norm regularized problem has also been applied in deep learning and neural networks problems [6, 7, 8].
@@ 2 Notation and Preliminaries
Let $\mathcal{X}$ and $\mathcal{Y}$ be two finite dimensional real Euclidean spaces. Given $\rho>0$, define the ball $\mathbb{B}_{\rho}(z):=\{z^{\prime}\in\mathcal{X}~{}|~{}\|z^{\prime}-z\|\leq\rho\}$, $\forall z\in\mathcal{X}$. Let $D\subseteq\mathcal{X}$ be a non-empty closed convex set. For any $z\in D$, the tangent cone (see [17, Definition 6.1]) of $D$ at $z$ is defined by ${\mathcal{T}}_{D}(z):=\{d\in\mathcal{X}~{}|~{}\exists z^{k}\rightarrow z$ with $z^{k}\in D$ and $\tau^{k}\downarrow 0$ s.t. $(z^{k}-z)/\tau^{k}\rightarrow d\}$. We use $\mathcal{N}_{D}(x):=\{d\in\mathcal{X}:\langle d,z-x\rangle\leq 0~{}\forall z\in D\}$ to denote the normal cone of $D$ at $x\in D$. Denote by $\delta_{D}(x)$ the indicator function over $D$, i.e., $\delta_{D}(x)=0$ if $x\in D$, and $\delta_{D}(x)=\infty$ if $x\notin D$. Define the support function of the set $D$ as $\sigma(y,D):=\sup_{x\in D}\langle x,y\rangle,y\in\mathcal{X}$. For a given $x\in\mathcal{X}$, define $\Pi_{D}(x):=\arg\min\{\|d-x\|~{}|~{}d\in D\}$ as the projection mapping onto $D$. Suppose the function $f:\mathcal{X}\rightarrow(-\infty,+\infty]$ is proper closed convex, denote by ${\rm dom}f:=\{x\mid f(x)<\infty\}$ the effective domain of $f$, by $f^{*}$ the conjugate of $f$ and by $\partial f$ the subdifferential of $f$. For more details, one can refer standard convex analysis [17]. For any convex cone $K\subset\mathcal{Y}$, denote by $K^{\circ}:=\{z^{\prime}\in\mathcal{Y}\mid\langle z^{\prime},z\rangle\leq 0,\forall z\in K\}$ the polar of $K$.
@@ 3 Variational Analysis of the Spectral Norm Function
In this section, we will provide a direct expression for the directional derivative of the proximal mapping of the spectral norm function. Then discuss the relationship between directional derivatives of the proximal mapping, the so called “sigma term” and the critical cones.
@@
    If $H\in\mathcal{C}_{\theta^{*}}(W,P)$ and $\varphi^{*}_{(W,H)}(P)=0$, then $H\in(\mathcal{C}_{\theta}(P,W))^{\circ}$.
@@
    If $H\in\mathcal{C}_{\theta}(P,W)$ and $\psi^{*}_{(P,H)}(W)=0$, then $H\in(\mathcal{C}_{\theta^{*}}(P,W))^{\circ}$.
@@
which means that $H\in(\mathcal{C}_{\theta}(P,W))^{\circ}$. Conversely, when $\tau^{\prime}$ in $\mathcal{C}_{\theta}(P,W)$ equals to 00, we can not get ${\rm trace}(\widetilde{H}_{\alpha\alpha})=0$, i.e., the reverse conclusion is not true. Then the proof of part (i) is completed.
To prove (ii), it is easy to see that $\psi^{*}_{(P,H)}(W)=0$ and $H\in\mathcal{C}_{\theta}(P,W)$ hold if and only if there exists $\tau^{\prime\prime}\in\Re$ such that
@@
which can be proved to be equivalent to $H\in(\mathcal{C}_{\theta^{*}}(W,P))^{\circ}$ by using the same method as in part (i). ∎
@@
Let $W\in\partial\theta(P)$ and $Q=P+W$ admit the SVD as in (3.1). Let the index sets $\alpha$, $\beta$, $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ and $c$ be defined as (3.2), (3.4) and (3.6). Given any $D\in\Re^{m\times n}$, denote $\widetilde{D}=U^{T}DV$. The function $\psi^{*}_{(P,D)}(W)$ is defined in (3.14). For any $D\in\Re^{m\times n}$, if $H={\rm Prox}^{\prime}_{\theta}(Q;H+D)$ holds, then $H\in\mathcal{C}_{\theta}(P,W)$ and $\langle H,D\rangle=-\psi^{*}_{(P,D)}(W)$.
@@
where $\Xi_{\alpha\alpha}$ is defined in (3.13). We first verify $H\in\mathcal{C}_{\theta}(P,W)$. The equality
@@
From part (i) of proposition 3.1, we obtain that $H\in\mathcal{C}_{\theta}(P,W)$.
@@
${\rm Prox}^{\prime}_{\theta}(Q;D)=0\Longleftrightarrow D\in(\mathcal{C}^{0}_{\theta}(P,W))^{\circ},$
where $\mathcal{C}^{0}_{\theta}(P,W)$ represents the critical cone $\mathcal{C}_{\theta}(P,W)$ in the case of $\tau=0$, i.e.,